# O-RAN xApps Conflict Management using Graph Convolutional Networks

### v1

@@ III-D Root Cause Analyst (RCA)
The RCA module is responsible for explaining the GAD module detections by providing the root causes (xApps) contributing to the detected conflicts (see Fig. 7). The detections from GAD and the whole constructed graph structure obtained from GSC are passed to the RCA module. The output is a subset of the graph of the detected conflicts.
Figure 7: The root cause analyst module steps.
The subset of the graph corresponding to each detection is further analyzed. In particular, it checks the source node with more than one incoming edge, as shown in Fig. 1. This means that the behavior of this node is being altered by more than one node. Then, it extracts the connected destination nodes as the root causes contributing to the detected conflict. The root causes are logged in a CSV file.
@@ IV Experiments and Results
This section presents the evaluation metrics, experimental setup, and performance metrics results of the proposed solution. The experimental setup outlines the environment, tools, and methodologies employed to assess the effectiveness of GRACE, while the results provide a comprehensive interpretation and analysis of its performance.
@@ IV-A Experimental Setup
We simulated and tested our model on an NVIDIA DGX station Intel(R) Xeon(R) CPU ES-2698 v4 @2.20GHz, 20core/40ht with 256GB of memory. The model is trained using the Adam optimizer with a learning rate of 0.01, the ReLU activation function, and a weight decay of 1e-4. The training process utilizes a batch size of 128, employing 5-fold cross-validation over 1,000 epochs. A two-layer GCN model is used to operates on the graph’s node features. It performs graph convolutions and aggregates information from the graph structure. Global mean pooling [26] is applied to return batch-wise graph-level-outputs by averaging node features across the node dimension.
@@ IV-B Experiment Results
We conducted a parametric study provided by TABLE I to systematically investigate how $\gamma$ value influences the model’s performance. The importance of this step lies in its ability to provide deeper insights into how $\gamma$ contributes to the model’s behavior. Understanding this correlation allows us to make better-informed judgments during the optimization phase.
@@
For the balanced dataset, $\gamma$ is set to zero in the focal loss function as the classes are equally distributed. All the classes contribute similarly to the loss function and the model is less likely to be biased toward the majority class. There is no need to emphasize the harder examples or penalize the easier ones. The model demonstrates significant performance on the balanced dataset, achieving high precision, recall, and F1 of 0.9829, 0.9818, and 0.9817, respectively.
@@
We evaluated different $\gamma$ values, ranging from 0.0 to 4.0, for the imbalanced datasets under investigation, as outlined in TABLE I. As the value of $\gamma$ increases, the performance metrics improve until $\gamma$ reaches 2.5 for the dataset with 40% conflict, where a high F1 of 0.9827 is achieved. Conversely, for the datasets with conflict percentages ranging from 30% to 10%, a $\gamma$ value of 3.5 yields superior performance. Specifically, the F1 for these datasets are 0.9864, 0.9924, and 0.9912, respectively. After this point, the metrics values decrease. This reduction in performance is expected, as excessively high values of $\gamma$ cause the model to overly focus on misclassified examples, potentially leading to overfitting. This, in turn, reduces the model’s ability to generalize to easier examples, resulting in a decrease in overall performance.
@@
The variation of the $\gamma$ parameter significantly influences the model’s performance. As $\gamma$ increases, the model places greater emphasis on hard-to-classify examples, thereby enhancing recall by improving its ability to identify difficult or underrepresented cases. However, this focus on challenging examples can lead to a reduction in precision, as the model may misclassify some simpler instances. Conversely, when $\gamma$ is set to a lower value, the model treats all examples with equal importance, resulting in a more balanced performance. However, this may limit the model’s capacity to adequately focus on difficult cases. Therefore, selecting the optimal $\gamma$ value is crucial for achieving a balance between precision and recall, ensuring the model performs effectively without overfocusing on any particular type of conflict.
The model exhibits high precision, recall, and F1 for the datasets under study, indicating its excellent performance in classifying data. The high precision indicates that when the model detects a positive class, it is correct most of the time, minimizing false positives. The high recall shows that the model is effective at identifying most of the true positive instances, reducing false negatives. The F1 demonstrates that the model achieves an outstanding balance between precision and recall since it balances these two metrics. These findings imply that the model is accurate and reliable, correctly classifying the conflicts while avoiding significant classification errors.
TABLE I: GCN Performance in Different Setups
Datasets Performance Metrics $\gamma$=0.0 $\gamma$=0.5 $\gamma$=1.0 $\gamma$=1.5 $\gamma$=2.0 $\gamma$=2.5 $\gamma$=3.0 $\gamma$=3.5 $\gamma$=4.0 40% conflict Prec 0.8892 0.9529 0.9592 0.9543 0.9755 0.9843 0.9546 0.9525 0.7242 Rec 0.8686 0.9200 0.9486 0.9429 0.9714 0.9829 0.9429 0.9257 0.8000 F1 0.8449 0.9139 0.9469 0.9392 0.9710 0.9827 0.9394 0.9183 0.7501 30% conflict Prec 0.7831 0.8850 0.9428 0.9709 0.9699 0.9778 0.9818 0.9881 0.9755 Rec 0.8267 0.8600 0.9267 0.9667 0.9600 0.9733 0.9800 0.9867 0.9733 F1 0.7993 0.8394 0.9261 0.9642 0.9579 0.9723 0.9785 0.9864 0.9722 20% conflict Prec 0.8526 0.8475 0.8925 0.9748 0.9805 0.9751 0.9469 0.9934 0.9047 Rec 0.9000 0.8846 0.9231 0.9692 0.9769 0.9692 0.9385 0.9923 0.9077 F1 0.8729 0.8460 0.9054 0.9656 0.9747 0.9658 0.9312 0.9924 0.9057 10% conflict Prec 0.8486 0.8200 0.8183 0.8716 0.8810 0.9226 0.9240 0.9942 0.9382 Rec 0.9035 0.9035 0.8947 0.9211 0.9386 0.9298 0.9474 0.9912 0.9298 F1 0.8716 0.8704 0.8509 0.8916 0.9089 0.9162 0.9321 0.9912 0.9200
@@
The training and validation losses plot for the 10% conflict dataset ($\gamma$ = 3.5) is shown in Fig. 8. The losses decrease throughout training. This indicates that the model is effectively learning from the data over time. Initially, both losses were high. This reflects the model’s early stage of learning, where it struggles to fit the data. As the training progresses, the loss decreases, indicating that the model is gradually learning to make better detections on the training data. Thus, the model improves its ability to fit the training data and gradually minimizes the error between its detections and the actual values. Similarly, the validation loss decreases, suggesting that the model is generalizing well and not overfitting to the training set. Both losses decrease at a similar rate and converge to a low value of less than 0.05. The trend of decreasing losses typically points to a well-optimized model, learning meaningful patterns and avoiding excessive complexity or bias.
Figure 8: GRACE training and validation loss per epoch on the 10% conflict dataset.
The confusion matrix (CM) of the proposed method testing on the 10% conflict dataset is elucidated in Fig. 9. CM provides a comprehensive overview of how well the model is performing by showing the number of the detected class and the true class. It allows us to identify if the model is biased toward the majority classes, especially in our case, where there are highly imbalanced datasets. The test set is composed of 114 instances. Out of the 114 instances, we have 108 normal, 1 direct conflict, 3 implicit conflicts (where 1 is misclassified as indirect conflict), and 2 indirect conflicts. By analyzing the figure, the diagonal elements of the matrix where the label and detected labels match are larger than the off-diagonal elements, which represent misclassifications. This indicates that the model can generalize well to new unseen data.
Figure 9: GRACE testing confusion matrix on the 10% conflict dataset.
@@
The conflict RCA report for the 10% conflict test dataset is illustrated in Fig. 10. The snapshot is filtered to show the detected conflict root causes. It indicates the detected label, the type of the conflict, the affected node, and the root causes (nodes and xApps). The affected node is the one whose behavior is altered by one or more nodes. The root causes nodes are the nodes that are concurrently modifying the same node. The root causes (xApps) are the xApps that are contributing to the detected conflict. The first line of the report shows a direct conflict occurrence (the detected label is 1) between two xApps (a4, and a7). In particular, these two applications control the same parameter p7. Thus, the root causes of this detected direct conflict are a4, and a7.
The experimental results validate the performance of our proposed model which is a great candidate to address highly imbalanced datasets. The model demonstrates strong generalization capabilities, making it well-suited for practical application in real-world scenarios.
Figure 10: Conflicts RCA report for the 10% conflict dataset.
V Conclusion
@@
In this work, we proposed a novel data-driven GCN-based method called GRACE for xApps conflict management and RCA in O-RAN. GRACE takes a binary-state dataset as input for the xApps, parameters, and KPIs in the network and constructs graph-structured data. Subsequently, it leverages GCN with a focal loss function to detect the occurrence of three types of conflicts (direct, indirect, and implicit). The learned graph structure and the detections are exploited to determine the root causes (xApps) that are contributing to the detected conflicts. The model demonstrates exceptional performance on highly imbalanced datasets. With conflicts ranging from 40% to 10%, the proposed model achieves a high F1 score greater than 98% on various evaluation metrics. Future work will focus on developing predictive maintenance with mitigation strategies based on the retrieved root causes to prevent performance degradation and reduce xApp conflicts across multi-vendor environments. Additionally, we plan to deploy and test the model in a real-life environment to further enhance its performance and applicability.

### v2

@@ III-D Root Cause Analyst (RCA)
The RCA module is responsible for explaining the GAP module predictions by providing the root causes (xApps) contributing to the predicted conflicts (see Fig. 7). The predictions from GAP and the whole constructed graph structure obtained from GSC are passed to the RCA module. The output is a subset of the graph of the predicted conflicts.
Figure 7: The root cause analyst (RCA) module.
The subset of the graph corresponding to each prediction is further analyzed. RCA checks the source node with more than one incoming edge, as shown in Fig. 1. This means that the behavior of this node is being altered by more than one node. Then, RCA extracts the connected destination nodes as the root causes contributing to the predicted conflict.
@@ IV Experiments and Results
This section presents the experimental setup and performance of the proposed solution. The experimental setup outlines the environment, tools, and methodologies employed to assess the effectiveness of GRAPHICA, while the results provide a comprehensive interpretation and analysis of its performance.
@@ IV-A Experimental Setup
We simulated and tested our model on an NVIDIA DGX station Intel(R) Xeon(R) CPU ES-2698 v4 @2.20GHz, 20core/40ht with 256GB of memory. The model is trained using the Adam optimizer with a learning rate 0.01, ReLU activation function, and a weight decay 1e-4. The training process utilizes a batch size of 128, employing 5-fold cross-validation over 2,000 epochs. A two-layer GCN model operates on the graph’s node features. It performs graph convolutions and aggregates information from the graph structure. Global mean pooling [25] is applied to return batch-wise graph-level-outputs by averaging node features across the node dimension.
@@ IV-B Experiment Results
We conducted a parametric study provided by Table I to demonstrate the model’s performance under different values of $\gamma$, where the imbalance may have a strong impact on the model’s effectiveness.
@@
For the balanced dataset, $\gamma$ is set to zero in the focal loss function as the classes are equally distributed. All the classes contribute similarly to the loss function, and the model is less likely to be biased toward the majority class. The model demonstrates remarkable performance on the balanced dataset, achieving high precision, recall, and F1 of 0.9829, 0.9818, and 0.9817, respectively.
@@
We evaluated the proposed model for more $\gamma$ values, spanning from 0.0 to 4.0, and computed the average performance across ten experiments. This analysis was performed on the imbalanced datasets under investigation, as shown in Table I. As the value of $\gamma$ increases, the performance metrics improve for all the datasets. The best performance is attained when $\gamma$ reaches 1.5 for the 40% conflict dataset and 2.0 for the 30%, 20%, and 10% conflict datasets, respectively. At these values, the model achieves an F1-score of 0.9930, 0.9946, 0.9908, and 0.9954, respectively. Afterward, the performance metrics begin to decrease. This reduction in performance is expected, as excessively high values of $\gamma$ cause the model to overly focus on misclassified examples, potentially leading to overfitting. This, in turn, reduces the model’s ability to generalize to easier examples, resulting in a decrease in overall performance.
@@
The variation of $\gamma$ significantly influences the model’s performance. As $\gamma$ increases, the model emphasizes hard-to-classify examples during the training phase, thereby enhancing recall by improving its ability to identify difficult or underrepresented cases. However, this focus on challenging examples can reduce precision, as the model may misclassify some simpler instances. Conversely, when $\gamma$ is set to a lower value, the model treats all examples with equal importance during training, resulting in a more balanced performance. However, this may limit the model’s capacity to focus on difficult cases adequately. Therefore, selecting the optimal $\gamma$ value is crucial for achieving a balance between precision and recall, ensuring the model performs effectively without overfocusing on any particular type of conflict. Overall, as shown in Table I, the model exhibits consistent precision, recall, and F1 across different test datasets with varying levels of imbalance.
TABLE I: Average Performance of GCN in different setups
Test Datasets Performance $\gamma$=0.0 $\gamma$=0.5 $\gamma$=1.0 $\gamma$=1.5 $\gamma$=2.0 $\gamma$=2.5 $\gamma$=3.0 $\gamma$=3.5 $\gamma$=4.0 40% conflict Prec 0.9849 0.9879 0.9900 0.9937 0.9923 0.9930 0.9923 0.9889 0.9892 Rec 0.9809 0.9861 0.9890 0.9931 0.9913 0.9919 0.9913 0.9873 0.9867 F1 0.9802 0.9859 0.9889 0.9930 0.9914 0.9918 0.9912 0.9871 0.9867 30% conflict Prec 0.9837 0.9884 0.9897 0.9929 0.9951 0.9915 0.9904 0.9906 0.9886 Rec 0.9804 0.9865 0.9878 0.9919 0.9946 0.9899 0.9878 0.9892 0.9872 F1 0.9795 0.9864 0.9876 0.9919 0.9946 0.9898 0.9875 0.9888 0.9869 20% conflict Prec 0.9648 0.9633 0.9792 0.9848 0.9939 0.9854 0.9834 0.9885 0.9867 Rec 0.9633 0.9641 0.9828 0.9836 0.9914 0.9820 0.9836 0.9828 0.9820 F1 0.9568 0.9586 0.9797 0.9824 0.9908 0.9818 0.9820 0.9814 0.9807 10% conflict Prec 0.9721 0.9760 0.9913 0.9954 0.9960 0.9871 0.9894 0.9766 0.9605 Rec 0.9816 0.9842 0.9798 0.9947 0.9956 0.9860 0.9895 0.9781 0.9702 F1 0.9757 0.9792 0.9798 0.9945 0.9954 0.9854 0.9878 0.9735 0.9630
@@
For further illustration, the training and validation losses across epochs for 10 experiments for the 10% conflict dataset ($\gamma$ = 2.0) are shown in Fig. 8, represented as shaded regions to show variability, with bold lines indicating the average training and validation trends. The losses decrease throughout training. This indicates that the model is effectively learning from the data over time. Initially, both losses were high. This reflects the model’s early stage of learning, where it struggles to fit the data. As the training progresses, the loss decreases, indicating that the model is gradually learning to make better predictions on the training data. Thus, the model improves its ability to fit the training data and gradually minimizes the error between its predictions and the actual values. Similarly, the validation loss decreases, suggesting that the model is generalizing well and not overfitting to the training set. Both losses decrease at a similar rate and converge to a low value of less than 0.05. The decreasing losses typically point to a well-optimized model, learning meaningful patterns and avoiding excessive complexity or bias.
Figure 8: Training and validation losses per epoch on the 10% conflict dataset.
The confusion matrix (CM) of the proposed method testing on the 10% conflict dataset ($\gamma$ = 2.0) is shown in Fig. 9. The test set consists of 114 data points. Among these, 97 are predicted as normal, 6 as direct conflicts, 3 as implicit conflicts, and 7 as indirect conflicts. In this highly imbalanced scenario, where the normal cases significantly outnumber the conflicts, only 1 implicit conflict was misclassified as an indirect one. This indicates that the model can generalize well in imbalanced data.
TABLE II: An example of conflict report by the RCA (10% conflict)
Predicted Label Conflict Type Affected Node Root Causes Nodes Root Causes xApps 1 Direct p12 a1, a3 a1, a3 3 Indirect p8 p1, p3 a1, a3 1 Direct p2 a4, a6 a4, a6 2 Implicit k10 p1, p9 a1, a9 1 Direct p10 a2, a5, a8 a2, a5, a8 3 Indirect p13 p7, p9, p10 a7, a9, a10 2 Implicit k5 p4, p2 a4, a2
@@
The conflict RCA report for the 10% conflict test dataset ($\gamma$ = 2.0) is illustrated in Table II. The normal data points in the table are filtered to emphasize the predicted conflicts. The report elucidates the predicted label, the type of the conflict, the affected node, and the root causes (nodes and xApps). The affected node is the one whose behavior is altered by one or more nodes. The root causes nodes are the nodes that are concurrently modifying the same node, and the root causes (xApps) are the applications contributing to the predicted conflict.
Figure 9: Confusion matrix on the 10% conflict test dataset.
As an example, the first line of the report indicates the occurrence of a direct conflict (predicted label = 1). The value of the parameter p12 is influenced by the xApps a1 and a3. Consequently, the root causes of this direct conflict are identified as the xApps a1 and a3. The last line of the report reveals an implicit conflict, with the predicted label being 2. In particular, the value of kpi k5 is affected by two parameters (p4 and p2), which are controlled by the xApps a4 and a2, respectively. Therefore, the root causes of this conflict are the xApps a4 and a2.
V Discussions
While the demonstrated performance is outstanding on the curated datasets, to the best of our knowledge, there is a lack of publicly available datasets or benchmarks for comparing the performance against other approaches to xApp conflict analysis. In [18, 17, 21], the proposed methods are tested on simulated synthetic data based on the Gaussian distribution function to generate KPIs for xApps. In [20] and [19], simulations are developed using network emulators such as the Colosseum wireless network emulator and “mobile-env” wireless communication environment in the Gymnasium framework, respectively. A key extension of the work is to enrich the scenarios and release, where applicable, a benchmark with open data from real or realistic O-RAN in 5G/beyond 5G. This would allow further validation and comparisons of different methods in practical scenarios.
Our results are still sufficiently good without benchmarking against others because the missed conflicts, although mostly implicit, are generally mild. The more severe conflicts are effectively flagged promptly. Our approach is particularly effective in practical applications, as it efficiently enables early prediction of rare conflict scenarios up to 10%, providing ample time for preventive measures to be put in place.
Compared to other works such as the one presented in [21], which claims an exceptionally high performance of 100%, our method is not inherently underperforming. Instead, it avoids reliance on potentially unrealistic benchmarks. Existing methods [18, 17, 21] are heavily distribution-dependent, often predicated on the assumption that the underlying data follows a Gaussian distribution, which is used to assign values to KPIs. However, in practical, real-world applications, KPIs frequently exhibit more complex behaviors and may not conform to Gaussian distributions. Our approach addresses this limitation by being distribution-independent, where KPIs are assigned binary state values based on their dynamic behavior. This design enables our methodology to generalize effectively across diverse KPI contexts without being constrained to specific distributions. Consequently, it can be argued that our method offers a superior level of adaptability and robustness for applications where distribution assumptions may not hold.
VI Conclusion
@@
In this work, we proposed a novel data-driven GCN-based method called GRAPHICA for xApps conflict management and RCA in O-RAN. Our approach takes a binary-state dataset as input and constructs graph-structured data. The binary state dataset encapsulates the dynamic variations in the values of the xApps subscriptions, controllable parameters, and KPIs within the network. Our Proposal leverages focal loss function to predict three types of conflicts (direct, indirect, and implicit). The RCA module then processes the learned graph structure and the predictions to determine the xApps being the source of the predicted conflict. The model demonstrates high performance on highly imbalanced datasets. With a presence of conflicts ranging from 40% to 10%, the proposed model architecture achieves an F1 score greater than 98% on various evaluation metrics for all the synthesized datasets. Future work will focus on developing predictive maintenance with mitigation strategies based on the retrieved root causes to prevent performance degradation and reduce xApp conflicts across multi-vendor environments. Additionally, we plan to deploy and test the model in a real-life environment to further enhance its performance and applicability.